OpenStudy (anonymous):

The formula for degrees Celsius is: C = (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.

OpenStudy (anonymous):

Part 1: c=f-32 +32 +32 f=c+32 Part 2: f=c+32 f=6+32 f=38

OpenStudy (anonymous):

thank you!(:

OpenStudy (anonymous):

No Problem! :)